OpenStudy (anonymous):

Which of the following compounds contains a greater percentage of water by mass, SnCl2 . 2H2O or NaC2H3O2 . 3H20?

OpenStudy (xishem):

Compound 1:\[1mol(SnCl_2)*\frac{189.62g(SnCl_2)}{1mol(SnCl_2)}=189.62g(SnCl_2)\]\[2mol(H_2O)*\frac{18.02g(H_2O)}{1mol(H_2O)}=36.04g(H_2O)\]\[Percent(H_2O)_{mass}=\frac{mass(H_2O)}{mass(SnCl_2·2H_2O)}\]\[Percent(H_2O)_{mass}=\frac{36.04g}{189.62g+36.04g}=0.1597=15.97percent\] Now for compound two, you would go through a similar process. Go ahead and try to do it for compound and see if you can come up with an answer.

OpenStudy (xishem):

By the way, for the second compound you should get... \[0.3972=39.72percent\]Therefore, the second compound contains a greater percentage of water by mass.